What are the concentrations of HSO4, SO42−, and H+ in a 0.37 M KHSO4 solution?
(Hint: H2SO4 is a strong acid; Ka for HSO4 = 1.3

×

10−2.)
Answer #1

We Have

Ka for HSO4- = 1.3 \times 10-2 = 0.013

[HSO4-] = 0.37 M

Now lets see the ICE table

HSO4-   \rightleftharpoons H+ + SO42-  

Initial 0.37    0 0

Change -x +x +x

Equilibrium 0.37-x x x

Equilibrium constant for above reaction is

Ka = [H+][SO42-]/[HSO4-] = (x)(x) /(0.37-x) = x2 / (0.37-x)

x2 / (0.37-x) = Ka = 1.3 \times 10-2 = 0.013

as x is too small than 0.37 , 0.37-x= 0.37

then

x2 /0.37 = 0.013

x2 = 0.00481

taking square roots on both the sides we get

x = 0.069 (or approx 0.07 M)

So

[H+] = x = 0.07 M

[SO42-] = x= 0.07 M

[HSO4-] = 0.37 -x = 0.37 - 0.07 = 0.30 M